OpenStudy (anonymous):

Are there any tricks for finding the derivative of a function containing sqr rt?

OpenStudy (anonymous):

One of the problems I'm working on is \[4\sqrt{x}-x\]

OpenStudy (anonymous):

can hrlp this to u

OpenStudy (lgbasallote):

indeed there is \[\huge \frac{d}{dx} \left(\sqrt{f(x)}\right) = \frac{f(x)}{2\sqrt{f(x)}}\]

hartnn (hartnn):

\(\sqrt{x}=x^{(1/2)}\\ then \quad use \quad (d/dx) x^n=nx^{n-1}\)

OpenStudy (anonymous):

I would write it out... sqrt(x) is x^(1/2) when you subtract 1 from 1/2 you can that easily... -1/2

OpenStudy (lgbasallote):

for example \[\huge \frac{d}{dx} \left(2\sqrt x\right)\] here, f(x) = x and f'(x) = 1 so you'll have \[2 \left(\frac{1}{2\sqrt x}\right)\] and simplifying \[\implies \frac{1}{\sqrt x}\]

OpenStudy (anonymous):

So for the problem I'm working on would it be 2x-1?

OpenStudy (lgbasallote):

not exactly

OpenStudy (lgbasallote):

\[\huge \frac{d}{dx} (4\sqrt x) \implies 4\left(\frac{1}{2\sqrt x}\right)\] do you get what i'm implying?

OpenStudy (anonymous):

Yes, but I'm still getting confused. I can't seem to get to the answer.

OpenStudy (anonymous):

Move the constant out from the equation and just find the derivative of sqrt(x) which can be written as: x^(1/2) Bring the 1/2 down in front and then you subtract 1 from 1/2... .5-1=-.5 We know that x^(-1/2) is negative, so it is in the denominator

OpenStudy (anonymous):

So \[\frac{ 4 }{ 2\sqrt{x} }-1\] would b the derivative?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Yes! Thanks for the help guys!